OpenStudy (anonymous):

AB= 5 4 B= 7 3 -2 3 2 1 find A. All matrix

OpenStudy (amistre64):

7a + 2b = 5 7c + 2d = -2 3a+1b = 4 3c+ 1d = 3 A = a b c d